When trying to find "x" and "y", how would i use the substitution method in a problem such as
x=y+3 and 2x-y=5 (same problem not two different ones)
1
Expert's answer
2011-02-14T05:12:43-0500
You have to substitute x=y+3 into the second equation:
2(y+3)-y = 5
2y + 6 - y = 5
y = 5-6 = -1
Thus x = -1+3 = 2.

Answer: (2,-1)